5x+4=6+10x can you help me
Answer by Nate(3500)

5x + 4 = 6 + 10x
5x - 5x + 4 = 6 + 10x - 5x
4 = 6 + 5x
4 - 6 = 6 - 6 + 5x
-2 = 5x
-2/5 = x